OpenStudy (biohazard9064):

Help Please Algebra 1

OpenStudy (biohazard9064):

Part A: Solve A = 9 over 2(x + 23) for x. (4 points) Part B: Determine the value of x when A = 108. (2 points) Part C: Solve –np – 90 > 30 for n. Show your work. (4 points)

OpenStudy (anonymous):

can you write the whole thing out =.= i don't understand

OpenStudy (biohazard9064):

thats the who thing

OpenStudy (biohazard9064):

here

OpenStudy (anonymous):

okay okay :3 got cha

OpenStudy (anonymous):

may i ask what chapter is this?

OpenStudy (biohazard9064):

module 2 2.10

OpenStudy (anonymous):

er.... what was that @.@

OpenStudy (biohazard9064):

?

OpenStudy (anonymous):

okay, you want me just give you the results or what

OpenStudy (biohazard9064):

yea

OpenStudy (anonymous):

A = -23 B = 1 C : –np – 90 > 30 -np > 120

OpenStudy (anonymous):

-n > 120/p n < -120/p

OpenStudy (anonymous):

okay so for part A : you should get like this: A= 9/2 times (x+23) A=9/2 times x + 207/2 A= 9x +207 A=x =23 Part B: you got A = 108 so, 108 = 9x+207 9x=-99 x=-11 Part C: -np-90>30 -np-60>0 -np>60 np<-60 n<-60 over P hope it help :)

OpenStudy (biohazard9064):

ok thanks

OpenStudy (anonymous):

ur welcome :)